# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### 12 Jan 2024 v8

@@ 3.2. Formulations of Machine Learning Tasks
where $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$, $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module (e.g., prompt model, adaptor, and LoRA) parameterized by $\mathcal{W}^{G}_{a}$, and $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for optimizing the global PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data $x^{C_{k}}$ and $y^{C_{k}}$ while keeping $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ frozen.
@@ 4. FTL-FM Taxonomy
In FTL-FM, we consider the following five research issues: (1) What to transfer; (2) How to transfer; (3) What to protect; (4) How to protect; (5) How to attack. Based on the three FTL-FM settings and the five research issues, we construct a taxonomy to categorize state-of-the-art FTL-FM works, as shown in Table 2.
@@
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., 2023c; Shen et al., $*{}^{*}$, WPE (Xu et al., $*{}^{*}$, PTB (Zhou et al., $*{}^{*}$ PPDS (Hou et al., 2023b; Yu et al., $*{}^{*}$ DP (Wang et al., $*{}^{*}$ SA (Fan et al., $*{}^{*}$ DP (Azam et al., 2023; Balunovic et al., $*{}^{*}$, FE (Gupta et al., $*{}^{*}$ HE (Jin et al., $*{}^{*}$ DP  (Xu et al., 2023a; Zhao et al., $*{}^{*}$, PEFT (Zhang et al., $*{}^{*}$, SA (Fan et al., $*{}^{*}$
@@
DP: Differential Privacy PPDG: Privacy-Preserving Data Generation MP: Model Trained via Proxy Data MDP: Masking-Differential Prompting
@@
Table 2. Taxonomy of existing FTL-FM works. This taxonomy only presents works focusing on the training phase.
5. FTL-FM: Foundation Models for Enhancing Domain Models
@@
5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly two categories of methods that adapt the knowledge of FMs to specific domains through data-level knowledge transfer: (1) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains; (2) a downstream client prompts tuning a closed-source FM (e.g., ChatGPT) by learning a local prompt generator based on domain-specific data and predictions generated by the FM. We summarize data-level knowledge transfer methods that fall into setting 1⃝ in Table 3222While we consider an FTL-FM approach should incorporate certain protection mechanisms to protect data or models, we still review highly relevant works that prioritize the development of knowledge transfer methods without considering data or model protection and include them in the Table to ensure thoroughness. (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. and elaborate on these methods in this subsection.
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
[…]
Mattern et al. (2022)
Kurakin et al. (2023)
Carranza et al. (2023)
BBT (Sun et al., 2022b, a) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + soft $P$
@@
Figure 6. Illustration of adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through data-level knowledge transfer. (a) illustrates that the knowledge of FM is transferred to downstream clients via privacy-preserving synthetic data sharing. (b) illustrates that a client can adapt knowledge of the FM to its local task via federated prompt learning.
In some proprietary domains, such as the audit department in a bank, data resources are typically limited for deep learning tasks. One potential solution to address data scarcity is leveraging FMs pre-trained on a large corpus of domain-specific data to generate synthetic training and testing data (e.g., LawGPT (Nguyen, 2023)). However, domain-specific data (e.g., financial and legal data) typically contain sensitive and private information. Thus, directly releasing FMs trained on these data presents potential privacy issues because sensitive information encoded in FMs can possibly be recovered by adversaries. In order to share synthetic data generated by FMs without compromising the privacy of training data, privacy-preserving data generation methods have been proposed (Carranza et al., 2023; Yue et al., 2023; Mattern et al., 2022; Kurakin et al., 2023). Essentially, these methods first fine-tune a pretrained public FM using differential privacy (DP) on private domain data and then exploit the DP-tuned FM to generate new data, which can be distributed to downstream clients to augment their training data (illustrated in Figure 6(a)). These methods correspond to stage 1 of objective 1⃝ formulated in Eq.(1) but with an additional step of generating data using fine-tuned FMs.
Alternatively, clients can adapt knowledge of the FM hosted by an FM server to their specific domains through federated prompt tuning. A typical setting of federated prompt tuning involves a client that learns a local model to generate prompts that can guide the FM toward the desired responses (illustrated in Figure 6(b)). For example, BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they cannot be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) and DSP (Li et al., 2023b) were proposed to learn a client side’s discrete (i.e., hard) prompt generator (DPG) for customizing FM’s responses. BDPL adopts a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize the DPG, while DSP optimizes the DPG through supervised fine-tuning and reinforcement learning. BBT, BBTv2, BlackVIP, BDPL, and DSP involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(7).
Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the federated prompt tuning with only one client to multiple clients, the workflow of which is illustrated in Figure  7. More specifically, Fed-BBPT and FedBPT involve two servers: an FM server, denoted as $S_{\text{FM}}$, and a federated server, denoted as $S_{\text{Fed}}$. The former hosts an FM to be prompted by clients, while the latter is for federated tuning clients’ local prompt generators. In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators based on local data and predictions generated by the FM, and the federated server aggregates clients’ local prompt generators to form the global prompt generator. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(7).
@@
Figure 7. Illustration of adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through data-level knowledge transfer. This figure illustrates that multiple clients adapt the knowledge of the FM hosted by an FM server to their local tasks via federated prompt tuning, and these clients also leverage the FedAvg algorithm or its variants with the help of the federated server to train their local domain models.
While these federated prompt tuning works can adapt knowledge of FMs to clients’ domain-specific tasks, they put the privacy of clients’ local data at high risk since they send prompts and input text directly to the FM server without any privacy protection. For applications where (training and inference) data contains sensitive and private information of users, privacy-preserving techniques such as DP-rewrite (Igamberdiev et al., 2022) should be applied to data transmitted between parties.
@@
Instead of leveraging prompt tuning, ZooPFL (Lu et al., 2023) adapts the knowledge of FM to downstream clients’ tasks by training a client-specific embedding and a semantic re-mapping module for each client. More specifically, each client appends the client-specific embedding to the embedding generated by a pre-trained auto-encoder, which aims to make the client’s local input consistent with the FM. Clients also collaboratively fine-tune encoders of their auto-encoders through FedAVG (McMahan et al., 2017) to share knowledge. The semantic re-mapping module aligns the logits output from the FM with the semantic space of clients’ local tasks. ZooPFL adheres to the objective formulated in Eq. (2) with $K>1$, $\eta=0$. The domain model $\mathcal{W}^{Ck}$ of each client $k$ in ZooPFL is the client-specific embedding and semantic re-mapping module.

### 07 Feb 2024 v10

@@ 3.2. Formulations of Machine Learning Tasks
where $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$ while $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module parameterized by $\mathcal{W}^{G}_{a}$ (e.g., prompt model, adaptor, and LoRA) and shared by all clients; $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for optimizing the global PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data $x^{C_{k}}$ and $y^{C_{k}}$ while $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is keeping frozen.
@@ 4. FTL-FM Taxonomy
We break down the general questions we raised in the introduction into the following five research issues: (1) What to transfer; (2) How to transfer; (3) What to protect; (4) How to protect; (5) How to attack. Based on the three FTL-FM settings and the five research issues, we construct a taxonomy to categorize state-of-the-art FTL-FM works, as shown in Table 2.
@@
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., 2023c; Shen et al., $*{}^{*}$, WPE (Xu et al., $*{}^{*}$, PTB (Zhou et al., $*{}^{*}$ PPDS (Hou et al., 2023b; Yu et al., $*{}^{*}$ DP (Wang et al., $*{}^{*}$ SA (Fan et al., $*{}^{*}$ DP (Azam et al., 2023; Balunovic et al., $*{}^{*}$, FE (Gupta et al., $*{}^{*}$ HE (Jin et al., $*{}^{*}$ DP  (Xu et al., 2023a; Zhao et al., $*{}^{*}$, PEFT (Zhang et al., $*{}^{*}$, SA (Fan et al., $*{}^{*}$ TEE (Huang et al., ${}^{*\dagger}$
@@
DP: Differential Privacy PPDG: Privacy-Preserving Data Generation MP: Model Trained via Proxy Data MDP: Masking-Differential Prompting TEE: Trusted Execution Environment
@@
Table 2. Taxonomy of existing FTL-FM works. This taxonomy only presents works focusing on the learning (e.g., training, fine-tuning, optimization) phase.
5. Foundation Models for Enhancing Domain Models
@@
Prompt Generator Optimization BBT (Sun et al., 2022b, a) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + soft $P$
@@
Discrete Prompt Optimization APE (Zhou et al., 2023b) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ Queries to initialize, evaluate, and generate $P$
DLN (Sordoni et al., 2023)
EvoPrompt (Hong et al., 2023) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ Queries to crossover, mutate, and evaluate $P$
PromptAgent (Wang et al., 2023a) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ Queries to retrieve error and generate error feedback; Queries to update $P$
LLM-generated Knowledge Transfer Fine-tune-CoT (Ho et al., 2023) Reasoning explanations from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
Prompting (Magister et al., 2023)
MT-COT (Li et al., 2022a)
D-SBS (Hsieh et al., 2023) Rationals for $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
Sci-CoT (Ma et al., 2023) Rationals for $X$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
PaD (Zhu et al., 2023b) Reasoning Programs from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
LaMini-LM (Wu et al., 2023c) Instructions from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
[…]
Mattern et al. (2022)
Kurakin et al. (2023)
Carranza et al. (2023)
5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly four categories of data-level knowledge transfer methods that adapt the knowledge of FMs to a client of a specific domain: (1) the client prompts tuning a closed-source FM (e.g., ChatGPT) by optimizing a local prompt generator based on its domain-specific data and predictions generated by the FM. (2) the client optimizes discrete prompts with the guidance of FMs. (3) the client fine-tunes its domain model with knowledge generated by FMs. (4) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains. We summarize data-level knowledge transfer methods that fall into setting 1⃝ in Table 3222While we consider an FTL-FM approach should incorporate certain protection mechanisms to protect data or models, we still review highly relevant works that prioritize the development of knowledge transfer methods without considering data or model protection and include them in the Table to ensure thoroughness. (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. and elaborate on these methods in this subsection.
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
Figure 6. Illustration of adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through prompt generator tuning. (a) illustrates that a client can adapt the knowledge of the server’s FM to its local task via training a local prompt generator. (b) extends the scenario of (a) to multiple clients. (b) illustrates that each client trains a local prompt generator supervised by predictions of the FM, and these clients also leverage the FedAvg algorithm or its variants with the help of the federated server to train their local domain models.
5.1.1. Prompt Generator Optimization
Alternatively, clients can adapt knowledge of the FM to their specific domains through federated prompt optimization. A typical setting of federated prompt optimization involves a client that learns a local prompt generator to generate prompts that can guide the FM toward the desired responses (illustrated in Figure 6(a)). For example, BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they cannot be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) and DSP (Li et al., 2023b) were proposed to learn a client side’s discrete (i.e., hard) prompt generator (DPG) for customizing FM’s responses. BDPL adopts a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize the DPG, while DSP optimizes the DPG through supervised fine-tuning and reinforcement learning. BBT, BBTv2, BlackVIP, BDPL, and DSP involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(7).
The aforementioned federated prompt tuning methods involve an FM server and only one client. Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the federated prompt tuning to multiple clients, the workflow of which is illustrated in Figure  6(b). More specifically, Fed-BBPT and FedBPT involve two servers: an FM server, denoted as $S_{\text{FM}}$, and a federated server, denoted as $S_{\text{Fed}}$. The former hosts an FM to be prompted by clients, while the latter is for federated tuning clients’ local prompt generators. In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators based on local data and predictions generated by the FM, and the federated server aggregates clients’ local prompt generators to form the global prompt generator. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(7).
@@ 5.1.1. Prompt Generator Optimization
ZooPFL (Lu et al., 2023) adapts the knowledge of FM to downstream clients’ tasks by training a client-specific embedding and a semantic re-mapping module for each client. More specifically, each client appends the client-specific embedding to the embedding generated by a pre-trained auto-encoder, which aims to make the client’s local input consistent with the FM. Clients also collaboratively fine-tune encoders of their auto-encoders through FedAVG (McMahan et al., 2017) to share knowledge. The semantic re-mapping module aligns the logits output from the FM with the semantic space of clients’ local tasks. ZooPFL adheres to the objective formulated in Eq. (2) with $K>1$, $\eta=0$. The domain model $\mathcal{W}^{Ck}$ of each client $k$ in ZooPFL is the client-specific embedding and semantic re-mapping module.
Figure 7. Illustration of federated discrete prompt tuning, LLM-generated knowledge transfer, and synthetic data transfer. (a) illustrates that a client adapts the knowledge of the FM to its local task by querying the FM to guide the discrete prompt optimization. The client may also turn to a local FM to provide guidance. (b) illustrates that a client queries an LLM to generate emergent ability knowledge and leverages this knowledge to fine-tune its local domain model. (c) illustrates that the knowledge of FM is transferred to downstream clients via privacy-preserving synthetic data sharing.
5.1.2. Discrete Prompt Optimization
Another line of research on federated prompt tuning is federated discrete prompt optimization (FedDPO), in which the client aims to find the optimal discrete prompt that maximizes the performance of the queried FM $\mathcal{W}^{S}_{\text{FM}}$ towards a score function $\mathcal{R}$ (e.g., accuracy) on certain validation data $\mathcal{D}_{\text{val}}$. The objective of the FedDPO is a variant of Eq.(7). We write this objective as follows:
(16) $\begin{split}P^{*}=\operatorname*{arg\,max}_{P}\mathcal{R}(f_{\mathcal{W}^{S}_{\text{FM}}}([X,P])|\mathcal{D}_{\text{val}})\end{split}$
where $X$ is the input, $P$ is the prompt and $[X,P]$ forms the query to the $f_{\mathcal{W}^{S}_{\text{FM}}}$ hosted by a server.
Figure 7(a) provides a high-level illustration of FedDPO. Generally, the discrete prompt optimization procedure involves three steps: evaluate existing prompts, select prompts with the best scores, and generate candidate prompts. The three steps iterate until the algorithm converges. During the iteration, the client keeps querying the server’s FM and possibly an auxiliary local FM to guide its prompt optimization. Various methods have been proposed to implement Eq.(16). In particular, APE (Zhou et al., 2023b) adopts Iterative Monte Carlo Search to explore the prompt search space. During the optimization, APE queries a black-box FM (i.e., GPT-3) to score candidate prompts and generate new prompts. DLN (Sordoni et al., 2023) extended APE and proposed Stacked LLMs to guide the prompt optimization. EvoPrompt (Guo et al., 2023) utilizes evolutionary algorithms to optimize prompts and leverage a black-box FM (i.e., GPT-3.5) to perform evolutionary operators such as crossover and mutation in the Genetic Algorithm. PromptAgent (Wang et al., 2023a) proposed an agent-based framework for optimizing prompts, aiming to produce expert-level task prompts via strategic planning and reflecting with error feedback during the optimization. Specifically, PromptAgent applies Monte Carlo Tree Search to orchestrate the prompt optimization: given a current prompt, the agent first queries an FM (i.e., GPT-3.5) to collect errors from the task dataset; next, it queries an FM (i.e., GPT-4) to provide error feedback, and then updates the prompt according to the feedback and transits to the next iteration.
While these federated prompt optimization methods (including prompt generator tuning and discrete prompt tuning) work well in adapting knowledge of FMs to clients’ domain-specific tasks, they put the privacy of clients’ local data at high risk since they send (soft or hard) prompts and input text directly to the FM server without any privacy protection. For applications where (training and inference) data contains sensitive and private information of users, privacy-preserving techniques such as DP-rewrite (Igamberdiev et al., 2022) should be applied to data transmitted between parties.
5.1.3. LLM-generated Knowledge Transfer
Large-scale FMs (i.e., LLMs) have demonstrated remarkable abilities, coined as emergent abilities (Wei et al., 2022), which encompass multiple captivating aspects, such as In-Context Learning (ICL), Chain-of-Thought (CoT), and Instruction Following (IF). LLM-generated Knowledge Distillation aims to efficiently fine-tune small LLMs for a client using knowledge generated by a powerful LLM. Figure 7(b) provides a high-level illustration of LLM-generated Knowledge Distillation.
Fine-tune-CoT (Ho et al., 2023), CoT Prompting (Magister et al., 2023) and MT-COT (Li et al., 2022a) prompt an LLM (e.g., GPT-3 or PaLM) to generate reasoning explanations and leverage these reasoning explanations to fine-tune a student model (e.g., T5). Distilling Step-by-Step (D-SBS) (Hsieh et al., 2023) prompts an LLM (e.g., PaLM) to generate labels along with rationales that justify the labels. D-SBS then leverages these rationales to fine-tune a student model (e.g., T5). In addition to rationales, D-SBS also exploits labels of the fine-tuning dataset as additional guidance to fine-tune the student model. Sci-CoT (Ma et al., 2023) proposed a similar approach as D-SBS, but it conducts the distillation through two small models: The first small model (i.e., Flan-T5-small) is fine-tuned by rationales generated by an LLM (i.e., GPT-3.5), aiming to generate rationales for questions. Subsequently, the second model (i.e., Flan-T5-small) is fine-tuned by the rationales generated by the first model, aiming to answer questions. PaD (Zhu et al., 2023b) queries GPT-3.5 to obtain reasoning programs for input questions and leverages the reasoning programs to fine-tune student models. PaD also leverages self-distillation as an additional supervision term during fine-tuning. To facilitate the deployment of language models in resource-constrained environments, LaMini-LM (Wu et al., 2023c) generates a large-scale offline distillation dataset comprising 2.58M instructions collected from various existing datasets and generated by ChatGPT (i.e., GPT-3.5), and fine-tune a collection of smaller language models (e.g., T5 and GPT2), which achieve comparable performance to Alpaca.
These methods involve only one client. They adheres to the objective formulated in Eq.(2) with $K=1$, $p^{C_{1}}=1$, $p^{S}=0$, and $\eta=0$. The specific loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ of training the client $1$’ domain model $\mathcal{W}^{C_{1}}$ using local data and knowledge transferred from server’s $\mathcal{W}^{S}_{\text{FM}}$ is formulated in Eq.(9). Currently, existing LLM-generated Knowledge Distillation methods primarily emphasize enhancing model performance and training efficiency, ignoring potential privacy concerns that may arise during the knowledge transfer process (as reported in Table 3). Nonetheless, privacy holds paramount importance in numerous real-life applications, particularly within the domains of finance and healthcare. Therefore, it is imperative to delve into the realm of privacy-preserving LLM-generated knowledge distillation, as it holds significant practical and research values.
5.1.4. Synthetic Data Transfer
In some proprietary domains, such as the audit department in a bank, data resources are typically limited for deep learning tasks. One potential solution to address data scarcity is leveraging FMs pre-trained on a large corpus of domain-specific data to generate synthetic training and testing data (e.g., LawGPT (Nguyen, 2023)). However, domain-specific data (e.g., financial and legal data) typically contain sensitive and private information. Thus, directly releasing FMs trained on these data presents potential privacy issues because sensitive information encoded in FMs can possibly be recovered by adversaries. In order to share synthetic data generated by FMs without compromising the privacy of training data, privacy-preserving data generation methods have been proposed (Carranza et al., 2023; Yue et al., 2023; Mattern et al., 2022; Kurakin et al., 2023). Essentially, these methods first fine-tune a pretrained public FM using differential privacy (DP) on private domain data and then exploit the DP-tuned FM to generate new data, which can be distributed to downstream clients to augment their training data (illustrated in Figure 7(c)). These methods correspond to stage 1 of objective 1⃝ formulated in Eq.(1) but with an additional step of generating data using fine-tuned FMs.